Two hundred thirty-five children go to camp Luther. There are 17 more girls than boys. How many are girls?

Found 2 solutions by math_helper, greenestamps:
Answer by math_helper(2461)


# girls = G
# boys = G - 17
G + (G-17) = 235
2G - 17 = 235
2G = 252
G = 126

Answer by greenestamps(13200)



An easy informal solution using logical reasoning -- which follows exactly the same path as the formal algebraic solution shown by the other tutor....

(1) Add 17 more boys (from a neighboring camp?) to make equal numbers of boys and girls and a total of 235+17 = 252.
(2) Half of that 252, which is 126, is the number of girls in the camp.